A rectangular field of area (81m*49m) is equal to the area of square field .find the length of one side of square field

1
Expert's answer
2021-12-27T13:28:57-0500

"\text{It is given that the area of a rectangle has length and width } 

\\
81 \text{ and } 49 \text{ respectively } 
\\
A_{rectangle}=L\times W=81 \times 49=A_{square}
\\
\text{Recall that } A_{square}=L^2=81 \times 49
\\
\implies L=\sqrt{81 \times 49}=9 \times7=63 \,(m)
\\
\therefore \text{The lenght of one side of square field is } 63 \, m."